-4y - (5y +6) = -7y + 3
Answer by funmath(2933)

-4y - (5y +6) = -7y + 3
-4y-5y-6=-7y+3
-9y-6=-7y+3
7y-9y-6=7y-7y+3
-2y-6=3
-2y-6+6=3+6
-2y=9
-2y/-2=9/-2
y=-9/2
:
Happy Calculating!!!